find the roots of the quadratic equation X^2– x + k =0 and the value of the constant

k given that one root of the equation is twice the other.


1
Expert's answer
2021-09-14T07:13:02-0400
"x^2-x+k=0"

By Viet Theorem


"x_1+x_2=-\dfrac{-1}{1}=1"

"x_1x_2=\dfrac{k}{1}=k"

Given "x_2=2x_1." Then


"x_1+2x_1=1"

"3x_1=1"

"x_1=\dfrac{1}{3}, x_2=\dfrac{2}{3}"

"k=\dfrac{1}{3}(\dfrac{2}{3})=\dfrac{2}{9}"

"\{\dfrac{1}{3}, \dfrac{2}{3}\}"


"k=\dfrac{2}{9}"